OpenStudy (anonymous):

The mean score for a standardized test is 1700 points. The results are normally distributed with a standard deviation of 75 points. If 10,000 students take the exam, how many would you expect to score between 1700 and 1775 points?

OpenStudy (e.cociuba):

(1700 - 1700) / 75 = 0 / 75 = 0 (1775 - 1700) / 75 = 75 / 75 = 1 P[0 < z < 1] = P[z = 1] - P[z = 0] P[z = 1] = 0.8413 P[z = 0] = 0.5 P[0 < z < 1] = 0.8413 - 0.5 = 0.3413 10000 * 0.3413 = 3413 people will score between 1700 and 1775

OpenStudy (anonymous):

so 3413 people?

OpenStudy (e.cociuba):

Yup :)

OpenStudy (anonymous):

i have three more if you would like help :)

OpenStudy (e.cociuba):

Sure, wats up?

OpenStudy (anonymous):

ill post them in other questiosn so you can get more medals :)

OpenStudy (e.cociuba):

k watev :)